OpenStudy (anonymous):

solve each equation system by the addition method -2x+3y=-7 3x+y=16

OpenStudy (anonymous):

-2x+3y=-7 ->eqn1 3x+y=16 ->eqn2 multiple equation 2 by 3

OpenStudy (anonymous):

i.e 3x*3 + y*3 = 16*3

OpenStudy (anonymous):

that is 9x + 3y = 48

OpenStudy (anonymous):

9x + 3y = 48 -> eqn 3

OpenStudy (anonymous):

add eqn 1 & eqn 3

OpenStudy (anonymous):

-2x+3y=-7 9x+3y=48 add

OpenStudy (anonymous):